OpenStudy (anonymous):

Solve the given initial-value problem by finding an appropriate integrating factor xdx+(yx^(2)+4y)dy=0, y(4)=0

OpenStudy (anonymous):

M=x, N=x^(2)y+4y My=0 Nx=2xy Nx-My/(M) =2xy-0/x=2y e^integral(2y)=e^(y^2)=integrating factor? Is this correct